OpenStudy (anonymous):

Given the polynomial function below, find F(-5). F(x) = x2 - 2x - 7 A. -42 B. 8 C. 28 D. -22

OpenStudy (alexandervonhumboldt2):

substitute -5 for x and solve

OpenStudy (alexandervonhumboldt2):

f(-5)=(-5)^2-2*(-5)-7=?

OpenStudy (anonymous):

is it D. -22

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

(-5)^2-2*(-5)-7=5^2+10-7=25+10-7=?

OpenStudy (alexandervonhumboldt2):

-*-=+ (-)^2=+^2

OpenStudy (anonymous):

wait is it A. -42

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (anonymous):

i horrible at algebra

OpenStudy (alexandervonhumboldt2):

well look: (-5)^2-2*(-5)-7=5^2+10-7=25+10-7=? can you simplify this?

OpenStudy (anonymous):

C. 28

OpenStudy (alexandervonhumboldt2):

yes